ABC manufactures door frames which it supplies mostly to house builders. The board of Jack is thinking of implementing a Balanced Scorecard (BSC) to help it manage its strategic performance. Detailed below are the financial performance measures currently used and the Board are looking for your help in the calculation of Residual Income (RI) and Return on Capital Employed (ROCE).

Y/E 31/12/2019

Y/E 31/12/2018

$000

$000

Revenue

510

543

Cost of sales

320

300

Gross profit

190

243

Other operating costs

99

90

Operating profit

91

153

Finance cost

26

20

Profit before tax

65

133

Tax

11

27

Profit after tax

54

106

Non-current assets

640

530

Current assets

380

325

Current liabilities

45

32

Non-current liabilities

190

100

WACC

9%

9%

Jack is also considering an investment to enhance production. This investment will shorten production time and reduce costs by the use of digital technology. The investment will have a useful life of four years, after which it will have a scrap value of $0. The financial details on this investment are listed below:

Investment:

Initial cost                                      $86.4 million

Net cash inflow                            $30 million per annum

Cost of capital                              9%

Depreciation policy                              Straight line

For information:

ROI is calculated using controllable profit. Controllable profit is calculated as net cash flows less depreciation.

For future planning purposes, ROI and Residual Income (RI) are calculated using the opening written down value of plant and licence assets in each year.

The Net Present Value (NPV) for the investment is positive.

                       

Required:

  1. Calculate the Return on Capital Employed (ROCE) and the Residual Income (RI) for Mars for the year end December 2019 period.  
  1. Prepare a summary schedule for the proposed investment for each of the four years ending 31st December 2020 to 31st December 2023 which shows the estimated value of:
    1. Controllable Profit per annum
    2. Residual income for years 1,2,3 and 4
    3. Return on Investment for years 1,2,3 and 4

  1. Identify one performance measure for each of the four categories of the Balanced Scorecard (BSC): Financial, Internal, Learning and Growth and Customers.

  1. Explain why two of the measures you identified in part c) above are appropriate for Jack.

  1. Discuss the nature of “cause and effect” in the BSC as it relates to your suggestions in part d) above.
Answer #1
(a) Return on Capital Employed
= (Earnings before Interest & Tax/Capital Employed) * 100
where, Earnings before Interest & Tax = Operating Profits
where, Capital Employed = Total Assets - Current Liabilities
Non-current assets     640.00
Current assets     380.00
Total assets 1,020.00
Less: Current liabilities      45.00
Capital employed     975.00
Earning before Interest & Tax 91
Return on Capital employed = (91/975)*100
= 0.0933 *100
= 9.33%
Residual Income = Operating Profits - (WACC * Average Capital Employed)
where,Operating Profits      91.00
WACC 9%
Average Capital Employed
= (Opening Capital Employed + Closing Capital Employed)/2
Opening
Capital
Employed
Closing
Capital
Employed
Non-current assets     530.00     640.00
Current assets     325.00     380.00
Total assets     855.00 1,020.00
Less: Current liabilities      32.00      45.00
Capital employed     823.00     975.00
Average Capital Employed = (823+975)/2
= 1798/2
= 899
Residual Income = 91 - ( 9% * 899)
= 91 - 80.91
= 10.09
(b)
Computation of Depreciation
Initial Cost      86.40
Less: Scrap Value           -  
Depreciable Value      86.40
Useful Life        4.00
Annual Depreciation = Depreciable Value/Useful Life
= 86.40/4
= 21.60
Controllable Profits per annum arising due the asset
= Net cash inflow - Depreciation
= 30 - 21.60
= 8.40
(i) Computation of Controllable Profits per annum
= Controllable Profits per annum before the usage of asset + Controllable
      Profits arising due to the asset
= 91 + 8.40
= 99.40
(ii) Residual Income = Operating Profits - (WACC * Capital Employed)
Year 1 Year 2 Year 3 Year 4
Operating Profits      99.40      99.40      99.40      99.40
Computation of Capital Employed
Year 1 Year 2 Year 3 Year 4
Non-current assets     640.00     640.00     640.00     640.00
Add:WDV of asset purchased      64.80      43.20      21.60           -  
Current assets     380.00     380.00     380.00     380.00
Total assets 1,084.80 1,063.20 1,041.60 1,020.00
Less: Current liabilities      45.00      45.00      45.00      45.00
Capital employed 1,039.80 1,018.20     996.60     975.00
Capital employed 1,039.80 1,018.20     996.60     975.00
WACC 9% 9% 9% 9%
WACC * Capital employed 93.58 91.64 89.69 87.75
Residual Income 5.82 7.76 9.71 11.65
(iiI) Return on investments = (Net Profit/Opening Capital Employed)*100
Computation of Net Income
Year 1 Year 2 Year 3 Year 4
Net income before purchase of
asset
     54.00      54.00      54.00      54.00
Add:Income arising due to purchase of asset        8.40        8.40        8.40        8.40
Net income      62.40      62.40      62.40      62.40
Year 1 Year 2 Year 3 Year 4
Non-current assets     640.00     640.00     640.00     640.00
Add:WDV of asset purchased      64.80      43.20      21.60           -  
Current assets     380.00     380.00     380.00     380.00
Total assets 1,084.80 1,063.20 1,041.60 1,020.00
Less: Current liabilities      45.00      45.00      45.00      45.00
Capital employed 1,039.80 1,018.20     996.60     975.00
Net income      62.40      62.40      62.40      62.40
Capital employed 1,039.80 1,018.20     996.60     975.00
Return on investments = (Net Profit/Opening Capital Employed)*100 6.00% 6.13% 6.26% 6.40%
(c.) Identification of performance measures for each category of BSC
Category Performance Measure
Finance Cash Flow
Increase in ROI
Internal Product Cycle Time
Unit Cost
Learning & Growth New Product Introduction
Customer On-time delivery
(d) The following two performance measures are appropriate for Jack
Increase in ROI (Finance) - The return on investments which was 5.54% in 2019 has
shown a consistent year on year increase after the purchase of the new asset.It has
increased from 6% in year 1 to 6.40% in year 4
Product Cycle Time (Internal) - The purchase of the new asset has resulted in shortening
of the production time which results in increase in quantity available to the company for
sale to the customers as well as reducing the unit cost of the product because the
fixed costs are spread over a higher quantity of products. This increase in efficiency
results in reduction of costs and increasing operating profits.
(e.)
The " cause and effect" model in balanced score card approach is used to describe as
to how a strategy will create change.The cause and effect model is captured by
identifying an objective and then creating a strategy map to achieve that objective.
Increase in ROI (Finance) - The cause for this performance measure is to increase the
returns on the investments made by the shareholders.The strategy adopted is to
increase the operating profits which would result in higher returns on investments.
Product Cycle Time (Internal) - The cause for this performance measure is achieving
customer satisfaction.The strategy adopted is to reduce the lead time for its customers
by reducing the product cycle time by purchasing a new asset which reduces the
product cycle by adopting digital technology